Please show work
Problems Solving Equations I 4 14 5 153 1. Solve for y:
Answer #1

Solution:

\small \frac{4y}{5}-\frac{14}{15}=\frac{y}{3}

Multiply both side of the equation by \small 15

\small \therefore 12y-14=5y

\small \therefore 12y-5y-14=5y-5y

\small \therefore 7y-14=0

\small \therefore 7y-14+14=0+14

\small \therefore 7y=14

\small \therefore y=2